OpenStudy (anonymous):

Determine the vertex of the given equation. y = 4x2 + 6x - 1 A.) (-3/4, -5/2) B.) (5/2, 103/6) C.) (-3/4, -13/4) D.) (5/6, 37/18)

OpenStudy (anteater):

There is more than one way you might find the vertex. Is there a particular method you like to use?

OpenStudy (anteater):

You could use x = -b/2a to find the x-coordinate, for example.

OpenStudy (anteater):

If you do that, what is the x-coordinate of your vertex?

OpenStudy (anonymous):

Let;s go with x = -b/2a. :)

OpenStudy (anteater):

Ok :)

OpenStudy (anteater):

Or you can put the equation in vertex form if you want to do a few extra steps. ;)

OpenStudy (anteater):

Please let me know what you find for the x-coordinate.

OpenStudy (anonymous):

The problem is, I have no idea what to do. My lesson doesn't explain this well. :(

OpenStudy (anteater):

Oh, ok! A quadratic equation is of the form: y = ax^2 + bx + c, where a, b, and c are the "coefficients" or numbers multiplied by your variables.

OpenStudy (anteater):

So, what are a, b, and c for your equation, then?

OpenStudy (anteater):

a = 4, b = ? , c = ?

OpenStudy (anonymous):

I'm sorry. :/

OpenStudy (anteater):

Your equation is y = 4x^2 + 6x - 1, so here a= 4, b= 6, and c = -1. Does that make sense so far?

OpenStudy (anteater):

That's ok, if it is new information to you , I don't mind adding extra explanation. :)

OpenStudy (anteater):

Just a moment, though, I have to finish ironing a shirt! BRB

OpenStudy (anteater):

Thanks for waiting!

OpenStudy (anonymous):

I don't mind waiting, as long as you help me. :)

OpenStudy (anteater):

So, since a = 4, b = 6, and c = -1 , and since the x-coordinate of the vertex is x = -b/2a, what is -b/2a for your problem?

OpenStudy (anonymous):

-6/16?

OpenStudy (anteater):

Close :) -6/8

OpenStudy (anteater):

Which simplifies to what?

OpenStudy (anonymous):

-3/4

OpenStudy (anteater):

Yes, so you can eliminate two of your choices, since they don't have an x-coordinate of -3/4

OpenStudy (anteater):

So B and D are tossed out

OpenStudy (anonymous):

Okay. :)

OpenStudy (anteater):

You have the x-coordinate of your vertex, x = -3/4 . How would you use that information to find the y-coordinate, do you think?

OpenStudy (anonymous):

Plug it into everywhere with an 'x'?

OpenStudy (anteater):

Yes :D

OpenStudy (anteater):

Put it in your equation for x, and then you can find y.

OpenStudy (anonymous):

The decimal for this number is -.75, right?

OpenStudy (anteater):

Yes. Please let me know what you get for y.

OpenStudy (anonymous):

Okay, thank you for the help. One moment!

OpenStudy (anteater):

Sure!

OpenStudy (anonymous):

What do I do with the -1?

OpenStudy (anteater):

Add it at the end. :)

OpenStudy (anteater):

So you have 4(-3/4)^2 + 6(-3/4) - 1 , or 4(-.75)^2 + 6(-.75) - 1

OpenStudy (anonymous):

I wouldn't do 4*4? I thought you multiplied the powers?

OpenStudy (anteater):

Yes, for the first term you would square the -3/4 first, and then multiply that result by 4. So that would give you (-3/4)(-3/4) = 9/16, and then 9/16 X 4/1 = 9/4 or 2.5

OpenStudy (anonymous):

This is what I done" 16 (-.75) + 6 (-.75) -1 -12 + 14.5 -1 16.5

OpenStudy (anteater):

You are close to being right. The only thing you did not do was to "square" the -.75 in the first term.

OpenStudy (anteater):

So, it should be 4(-.75)^2 + 6(-.75) - 1 or 4(-.75)(-.75) + 6(-.75) - 1

OpenStudy (anteater):

Since you are not squaring the coefficient (4), but the "x" (-.75)

OpenStudy (anonymous):

Okokok. I felt like I was messing up. :p

OpenStudy (anteater):

That's ok :) So if you simplify 4(-.75)(-.75) + 6(-.75) - 1 , that will be our y

OpenStudy (anteater):

And it should match one of the answers. ;D

OpenStudy (anteater):

Please let me know what you get, and I will do the calculation as well.

OpenStudy (anonymous):

What is 9/16 in decimal form?

OpenStudy (anteater):

Let's see if you divide 9 by 16, you get 0.5625

OpenStudy (anonymous):

That's what I got, I was making sure. :p

OpenStudy (anteater):

Ok, good :D

OpenStudy (anteater):

So then, y = ?

OpenStudy (anonymous):

I got -3.25

OpenStudy (anteater):

Good :D

OpenStudy (anteater):

And would -3.25 be -13/4 or -5/2?

OpenStudy (anonymous):

-5/2?

OpenStudy (anteater):

Do you have a calculator handy?

OpenStudy (anonymous):

I do. :p

OpenStudy (anteater):

You can think of a fraction as being a division (numerator divided by denominator)

OpenStudy (anteater):

So, divide -13 by 4 ; and divide -5 by 2. Which one gives you the answer -3.25?

OpenStudy (anonymous):

-13/4 :)

OpenStudy (anteater):

Yes :D

OpenStudy (anteater):

So ... which choice?

OpenStudy (anonymous):

C. :)

OpenStudy (anteater):

Yes, C is our perpetrator. ;)

OpenStudy (anteater):

You chose it from the lineup ;)

OpenStudy (anteater):

I hope this was helpful!

OpenStudy (anonymous):

That made me laugh! xD

OpenStudy (anteater):

Good! Math should be approached with a sense of humor, I think :D

OpenStudy (anteater):

Keeps you from going nuts while doing it ... I hope!

OpenStudy (anteater):

Have a very good morning, Ms. Alyjoan! I must now hand off the computer. Family computer, so I have to share. :)

OpenStudy (anteater):

Bye! :)

OpenStudy (anonymous):

Actually, are you busy, in case I need help again?

OpenStudy (anteater):

I will be available this evening, if that works?

OpenStudy (anteater):

And if you have questions, I will be happy to explain the steps a little better then. :)

OpenStudy (anonymous):

Tha'll be fine. :) Thank you!

OpenStudy (anteater):

What time zone, though? I am on the west coast, so Pacific time.

OpenStudy (anteater):

In the evening Pacific time. :)